OpenStudy (anonymous):

A sample of gas has a pressure of 3.2 atmospheres at a temperature of 301 Kelvin. What would be the pressure of the same sample of gas when the temperature is decreased to 285 Kelvin? Show all of the work used to solve this problem.

OpenStudy (anonymous):

3.2 for P1

thomaster (thomaster):

\(\Large\sf \dfrac{P_1}{T_1}=\dfrac{P_2}{T_2}\) P= pressure T= temperature which 3 of these 4 can you fill in?

OpenStudy (anonymous):

301 for T1

thomaster (thomaster):

Yes thats correct

OpenStudy (anonymous):

285 for T2

OpenStudy (anonymous):

and I guess were trying to figure out the pressure it would be at that temp.

thomaster (thomaster):

Now do you know how to rework this formula to get P2 in front?

OpenStudy (anonymous):

no :/

thomaster (thomaster):

\(\Large\sf \dfrac{P_1}{T_1}=\dfrac{P_2}{T_2}\) what happens when you multiply both sides by T2?

OpenStudy (anonymous):

i have no idea honestly

OpenStudy (anonymous):

I don't understand how to get it by itself

thomaster (thomaster):

Alright lets pretend P2 = 1 and T2 = 4 so you'll have \(\dfrac{1}{4}\) What happens when you multiply 1/4 by 4?

OpenStudy (anonymous):

you have 1

thomaster (thomaster):

yes, when you multiply a fraction by it's denominator (the bottom one), the answer will be the numerator (the upper one) So \(\large\sf \dfrac{P_2}{T_2}*T_2=P_2\)

OpenStudy (anonymous):

oh ok

thomaster (thomaster):

Now we need to multiply the fraction in front of the equal sign too You will get \(\large\sf \dfrac{P_1}{T_1}*T_2\) because you can't cancel anything out here. Now we have reworked the formula to be: \(\large\sf P_2=\dfrac{P_1}{T_1}*T_2\) Just plug in the numbers and you'll have your answer :)

OpenStudy (anonymous):

oh okay let me try it (:

OpenStudy (anonymous):

3.02?

OpenStudy (anonymous):

3.03*

thomaster (thomaster):

Yes that's correct, great job! The pressure at 285 kelvin = 3.03 atm

OpenStudy (anonymous):

ah thanks only 9 more till my assignments done haha. i'd appreciate more help but I don't wanna annoy you

thomaster (thomaster):

I'm here to help, so you won't annoy me with any question :) First try it yourself and when you're having trouble with the question, you can ask for help here.

OpenStudy (anonymous):

okay i'll just write the ones i know I can't do

thomaster (thomaster):

Don't forget to post it in a new questions :)